40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5463     Entry time: Mon Sep 19 16:20:35 2011
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: AS55 whitening gain decreased 

The gain of whitening filters on AS55 was decreased from 21 dB to 0 dB for the Y arm locking.

 

- - (Background)- -

Since the modulation depths became bigger from the past (#5462), the PDH signal from Y arm was saturated in the path of AS55.

Due to the saturation the lock of the Y arm became quite difficult so I decreased the gain of of the whitening filter from 21 dB to 0 dB.

In this condition, a required gain in C1:LSC-YARM_GAIN is about -0.3, which is 10 times bigger from the default number.

For the MICH locking tonight, it may need to be back to a big gain.

ELOG V3.1.3-